2x - 4x + 3 = 0
-3 + 4x - 2x = 0
in change

Answer by edjones(8007)

2x - 4x + 3 = 0
-2x+3=0
-2x=-3
x=3/2
-3 + 4x - 2x = 0
-3+2x=0
2x=3
x=3/2
.
Ed